A bear runs at John with a velocity of 12m/s. John stops the bear by pushing on it for 2 seconds in a motion opposing the bear’s. If the bear weighs 200kg, what average force is needed to stop the bear?

Answer: -1200N

The answer is given, but I would like to know how to do this problem.
Answer #1

Solution)

Given,

Initial velocity, u=12 m/s

Time, t=2 sec

Mass, m=200 kg

Final velocity, v=0

We know,

Average Force, F=m(v-u)/t

Substitute values,

F=200*(0-12)/2= -1200 N (Ans)

===========